# Multi-fidelity Gaussian process surrogate modeling for regression problems in physics

### v1

@@ 2.1 Gaussian process regression
After we observe the training data, the posterior consists of a narrowed distribution of functions, initially defined by the prior distribution as observed in Figure 1. The inference is done then using posterior mean and posterior covariance,
@@
Having posterior mean and covariance, we can generate predictions for unobserved points.
Figure 1: A kernel function defines the family of the functions with which we are approximating the target function. The first picture shows a prior distribution defined by a kernel. In the posterior distribution in the second plot, newly observed training data points, depicted here in black, narrow the space of potential functions from the initial family.
@@ 2.2 Kernels
$K$ denotes a matrix built with the help of a kernel, which is a function used to define the covariances between the data points. This covariance is usually interpreted as a measure of similarity between data points, and thus, points $x$ and $x^{\prime}$ that are considered similar in the input space will have target values $y$ and $y^{\prime}$ that are also similar. The covariance also encodes certain properties of the functions we want to predict, for example, periodicity [28]. Choosing the right kernel is crucial in the modeling process with GP.
A valid covariance function must be positive semidefinite [21]. The product and sum of two valid covariance functions generate a valid covariance function. Using this property, one can generate more sophisticated covariance functions. There are multiple options for covariance functions like squared-exponential, exponential, trigonometric, etc. The task of choosing the right kernel is complicated. There is no standardized algorithm to do so except for the random search and the researcher’s intuition regarding modeling. However, some authors are making promising progress by creating interesting heuristics and algorithms [29, 30, 31, 32]. In this work, we only work with squared-exponential covariance functions.
@@ 2.3 Calibration
As shown in equation 6, hyperparameter optimization for Gaussian process kernels is usually performed by maximizing the marginal likelihood. This often leads to an acceptable mean squared error [21], but the posterior variance tends to be lower than the variance of the actual distribution [34, 35]. This problem requires an additional post-processing step known as calibration. There exists a variety of calibration techniques, but in all cases, they aim to adjust the predicted variance towards the actual variance in the data. Calibration is an established concept in the classification context, but in the case of regression, it is relatively new [36]. In general, calibration can be split into three categories: quantile, distribution, and variance calibration.
@@
for the set of data ${(x_{n},y_{n})}^{N}_{n=1}$, where $F_{n}$ stands for predicted cumulative distribution at $x_{n}$ and $F^{-1}_{n}$ the corresponding quantile function. Intuitively, $X\%$ of the confidence interval should capture the ground truth values in $X\%$ predictions [37, 38]. Quantile calibration is considered a global calibration, as it concentrates only on the marginal distribution without considering a calibration of distribution around each exact prediction.
Distribution calibration resembles the calibration approach in classification tasks because it is also focused on local calibration. In the classification case, instances are grouped by their predicted probability and they are considered calibrated only if each of that group matches the indicated probability. Assume $X$ as input variables, $Y$ as a target variable, and the regression model $f:\mathbb{X}\rightarrow\mathbb{S}_{\mathbb{Y}}$ or intuitively, the regression model is mapping from input features into a space of distributions and for each test instance predicted a particular distribution. Denoting $s$ as an arbitrary distribution predicted by the regression model, the regression model $f$ is distribution calibrated if
@@
Different methods approach uncertainty calibration tasks in different ways, therefore it is beneficial to test several of them on given data.
@@ 3.1.1 Linear Auto-regressive model
$=\mathbb{E}\left[\left(\rho u_{l}+u_{\delta}\right)\left(\rho u_{l}^{\prime})+u_{\delta}^{\prime}\right)\right]=\rho^{2}\mathbb{E}[u_{l}u_{l}^{\prime}]+\mathbb{E}[u_{\delta}u_{\delta}^{\prime}]=\rho^{2}K_{l}+K_{\delta}.$
@@
Using the maximum likelihood method, one can calculate the value of $\rho$ and the kernel hyperparameters. We can also extend the joint distribution in equation 12 to incorporate more than two fidelities that will lead to covariance matrices with sparse blocks. Le Gratiet [23] took advantage of this block structure and decreased the complexity of the method from $\mathcal{O}\left((\sum_{l=1}^{L}N_{l})^{3}\right)$ to $\sum_{l=1}^{L}\mathcal{O}(N_{l}^{3})$. He also improved the performance using polynomial terms for scaling ($\rho$) instead of a constant term.
@@ 3.1.2 Non-linear auto-regressive models
where $g:\mathbb{R}^{d+1}\rightarrow\mathbb{R}$ is a function representing the non-linear transformation applied on the low-fidelity function to obtain the high-fidelity model. When attempting to model both the low-fidelity and high-fidelity functions using GPs, writing the joint distribution of the surrogates, as presented in equation 12, becomes analytically challenging for a general transformation $g$. To overcome this difficulty, a common assumption in studies such as [25, 24] is that the low-fidelity function can be evaluated at zero cost. This eliminates the need for a surrogate for the low-fidelity function. Consequently, the non-linear transformation depicted in equation 13 transforms into creating a surrogate in a $d+1$ dimensional space. In our work, we delve into the modeling of $g$ using a Gaussian Process (GP). This approach enables us to effectively navigate the challenges associated with the joint distribution of surrogates, providing a viable solution for incorporating non-linear transformations within the multi-fidelity surrogate modeling framework.
@@
Figure 2 illustrates the relationship between $f_{l}$ and $f_{h}$. Modeling the transformation $g$ as described in equation 13 is synonymous with learning the manifold depicted in Figure 2. Crucially, this manifold lacks oscillations, signifying that learning it necessitates relatively fewer high-fidelity function evaluations. This observation underscores the advantage of multi-fidelity modeling in this example, as it allows for a learning process by leveraging the complicated feature to be captured in the low-fidelity function, thereby reducing the requirement of high-fidelity function evaluations.
@@
Here, $\theta_{\rho}$, $\theta_{f}$, and $\theta_{\delta}$ denote the kernel hyperparameters, which are determined by maximizing the likelihood, as discussed in the previous section. This kernel structure closely mirrors the Auto-Regressive Model (AR1) formulation described in equation 10, assigning each kernel to an individual part of the transformation. Specifically, $k_{\rho}$ models the scaling term, $k_{f}$ is responsible for transforming the low-fidelity model, and $k_{\delta}$ handles the additive correction term. The formulation above gives the method its name: Non-linear Auto-Regressive Gaussian Process (NARGP).
@@
Substituting equation 13 into equation 15, we obtain
@@
The Taylor expansion also motivates us to incorporate the derivative of low-fidelity in equation 13. Lee et al. [24] suggest writing the non-linear transformation as
@@
In many legacy simulators, the gradients are unavailable [40]. In cases where gradients are unavailable, one can resort to finite difference techniques to approximate the derivative. Lee et al. [24] suggest replacing the derivative in equation 16 with the delay term to accommodate such scenarios. This delay term evaluates the low-fidelity function after a small time step $\tau\in\mathbb{R}$. For a higher-dimensional input variable, we add a delay along each dimension, defining $\tau_{i}$ as a vector of zeros with $\tau$ at the $i^{\textrm{th}}$ position. The transformed formulation is now expressed as
@@
This method is called the Gaussian Process with Delay Fusion (GPDF) proposed in [24]. Additionally, we propose an enhancement to this method by modifying the kernel to mimic the Taylor expansion using a structured composite kernel similar to equation 14
@@
Figure 3: Flowchart showing the training and prediction steps of NARGP surrogate for a three-fidelity scenario. Training the NARGP model for a particular level involves evaluating the model of that fidelity level and the previous fidelity level. We draw samples to make predictions for a particular level, which involves drawing samples from the surrogate of the previous level and feeding them as input to the surrogate of that level. We can then use the samples to obtain the required statistical moments of the predictions by marginalizing the samples.
@@ 3.1.3 Deep Gaussian Processes
We model the transformation similar to equation 13 with one difference. We replace the actual function evaluation of the previous layer with the previous GP layer
@@
Figure 4: Flowchart showing the training and prediction steps of Deep Gaussian process surrogate for three fidelity scenarios. Training the DGP model for a particular level involves evaluating the model of that fidelity level and the surrogate of the previous fidelity level, which further involves evaluating the one-level lower surrogate. This creates a cascading effect until we reach the lowest fidelity surrogate. The prediction process is similar to the full GP, which involves drawing and marginalizing samples.
@@
One can combine all the layers by multiplying the posterior from individual layers from equation 21 to obtain the joint posterior distribution for all layer $p(\mathbf{y},\{\mathbf{g}_{\ell}^{t},u_{\ell}\})$. Analytical inference from the posterior is intractable. To approximate the distribution, we assume that the prior $q(\mathbf{u}_{\ell})$ of $\mathbf{u}_{l}$ is a Gaussian distribution with mean $\tilde{\mathbf{m}}_{\ell}$ and variance $\tilde{\mathbf{S}}_{\ell}$. Now, the approximated DGP prior is
@@
If $\mathbf{u}_{\ell}$ is marginalized out of the equation 22, then the distribution becomes a Gaussian distribution with the following mean and variance:
@@
We can use the before approximating the posterior at each layer and then extend this to all the layers of DGP to obtain the approximate posterior distribution $q(\mathbf{y},\{\mathbf{g}_{\ell}^{\ell},\mathbf{u}_{\ell}\}_{\ell=1}^{L})$. We want to approximate the exact posterior distribution in equation 21 by an approximate distribution $q(\mathbf{y},\{\mathbf{g}_{\ell}^{\ell},\mathbf{u}_{\ell}\}_{\ell=1}^{L})$. We train the DGP by minimizing the Kullback-Leibler (KL) divergence $\textrm{KL}[q||p]$ between the variational posterior distribution $q$ and the true posterior distribution $p$. An exact evaluation of the KL divergence is not feasible. So, we convert the minimization of the KL divergence to the maximization of Evidence Lower Bound (ELBO) ($\mathcal{L}$) whose evaluation is feasible. ELBO is evaluated as
@@
The prediction process in Deep Gaussian Processes (DGP) follows a recursive transfer of results from the previous layer into the next layer. We draw samples at each layer from the approximate distribution defined by equation 22. We fuse the prediction parameter $X^{*}$ in intermediate layers with the sample drawn from the previous layer. We then use the fused parameter to draw a sample from that layer. We repeat this process until we have $S$ samples for each layer. We use the collected samples to compute the posterior mean and variance through a Monte Carlo approximation. Figure 4(b) visually illustrates these prediction steps for three fidelity models, depicting the recursive transfer and sampling process at each layer.
One can improve the surrogate by choosing a more structured kernel. Cutajar et al. [26] suggest using the kernel defined in equation 14. We call this method the Non-linear Auto-Regressive Deep Gaussian Process (NARDGP). We suggest using the delay term to further add useful information. We call the method Deep Gaussian Process with Delay Fusion (DGPDF). We further improve upon the suggested method using a structured kernel as mentioned in equation 18. We named the resulting method Deep Gaussian Process with Delay Fusion and Composite kernel (DGPDFC).
@@ 3.2 Adaptive model improvement
Upon simplification, it can be shown that for GPs, the gain in information is directly proportional to the logarithm of the posterior predicted standard deviation. Thus, the optimization problem stated in equation 27 can be converted to
@@
There are certain drawbacks to this method. Firstly, this method adds one point at a time, limiting its overall speed. Secondly, the suggested points are always towards the boundary in high-dimensional space. Moreover, there are many other methods suggested in the literature that rely on other metrics to sequentially improve the Gaussian Process surrogate [49, 50]. We leave the discussion and effects of those algorithms on multi-fidelity GP as future works.

### v2

@@ 2.1 Gaussian process regression
After we observe the training data, the posterior consists of a narrowed distribution of functions, illustrated in Figure 1(b), initially defined by the prior distribution as observed in Figure 1(a). The inference is done then using posterior mean and posterior covariance defined as
@@
(a) Visualization of Gaussian process prior.
(b) Visualization of Gaussian process posterior.
Figure 1: A kernel function defines the family of the functions with which we are approximating the target function. In this example, we use squared-exponential or RBF kernel and visualize a) a prior distribution over the functions. b) In the posterior distribution in the second plot, newly observed training data points, depicted here in black, narrow the space of potential functions from the initial family, defined by the RBF kernel.
@@ 2.2 Kernels
$K$ denotes a matrix built with the help of a kernel, which is a function used to define the covariances between the data points. This covariance is usually interpreted as a measure of similarity between data points, and thus, points $x$ and $x^{\prime}$ that are considered similar in the input space will have target values $y$ and $y^{\prime}$ that are also similar.
A valid covariance function must be positive semidefinite [21]. The product and sum of two valid covariance functions generate a valid covariance function. Using this property, one can generate more sophisticated covariance functions.
There are multiple options for covariance functions like squared-exponential, exponential, trigonometric, etc. The task of choosing the right kernel is complicated. There is no standardized algorithm to do so except for the random search and the researcher’s intuition regarding modeling. However, some authors are making promising progress by creating interesting heuristics and algorithms [28, 29, 30, 31]. In this work, we only work with squared-exponential covariance functions, because it yields smooth, i.e. mean-square differentiable, sample paths from the resulting GP, while others, like exponential kernel, do not [21].
@@ 2.3 Calibration
As shown in equation (6), hyperparameter optimization for Gaussian process kernels is usually performed by maximizing the marginal likelihood. This often leads to an acceptable mean squared error [21], but the posterior variance tends to be lower than the variance of the actual distribution [33, 34]. This problem requires an additional post-processing step known as calibration. There are various calibration techniques, but in all cases, they aim to adjust the predicted variance towards the actual variance in the data. Calibration is an established concept in the classification context, but in the case of regression, it is relatively new [35]. In general, calibration can be split into three categories: quantile, distribution, and variance calibration.
@@
for the set of data ${(x_{n},y_{n})}^{N}_{n=1}$, where $F_{n}$ stands for predicted cumulative distribution at $x_{n}$ and $F^{-1}_{n}$ the corresponding quantile function. Intuitively, $X\%$ of the confidence interval should capture the ground truth values in $X\%$ predictions [36, 37]. Quantile calibration is considered a global calibration, as it concentrates only on the marginal distribution without considering a distribution calibration around each exact prediction.
Distribution calibration resembles the calibration approach in classification tasks because it is also focused on local calibration. In the classification case, instances are grouped by their predicted probability and considered calibrated only if each group matches the indicated probability. Assume $X$ as input variables, $Y$ as a target variable, and the regression model $f:\mathbb{X}\rightarrow\mathbb{S}_{\mathbb{Y}}$ or intuitively, the regression model is mapping from input features into a space of distributions and for each test instance predicted a particular distribution. Denoting $s$ as an arbitrary distribution predicted by the regression model, the regression model $f$ is distribution calibrated if
@@
The first method performs a global calibration, while the last two perform a local one. Global calibration is a better choice in case we need only to rescale the marginalized distribution, but in the case when variance should be rescaled only in certain places, it is better to use local calibration. Knowing a priori the required type of rescaling is hard for real-world data, so if we do not have any prior assumptions on the real variance, we must empirically find the most suited calibration method.
@@ 3.1.1 Linear Auto-regressive model
$=\mathbb{E}\left[\left(\rho u_{l}+u_{\delta}\right)\left(\rho u_{l}^{\prime}\right)+u_{\delta}^{\prime}\right]=\rho^{2}\mathbb{E}[u_{l}u_{l}^{\prime}]+\mathbb{E}[u_{\delta}u_{\delta}^{\prime}]=\rho^{2}K_{l}+K_{\delta}.$
@@
Using the maximum likelihood method, one can calculate the value of $\rho$ and the kernel hyperparameters. We can also extend the joint distribution in equation (12) to incorporate more than two fidelities that will lead to covariance matrices with sparse blocks. Le Gratiet [23] took advantage of this block structure and decreased the complexity of the method from $\mathcal{O}\left((\sum_{l=1}^{L}N_{l})^{3}\right)$ to $\sum_{l=1}^{L}\mathcal{O}(N_{l}^{3})$. He also improved the performance using polynomial terms for scaling ($\rho$) instead of a constant term.
@@ 3.1.2 Non-linear auto-regressive models
where $g:\mathbb{R}^{d+1}\rightarrow\mathbb{R}$ is a function representing the non-linear transformation applied on the low-fidelity function to obtain the high-fidelity model. When attempting to model both the low-fidelity and high-fidelity functions using GPs, writing the joint distribution of the surrogates, as presented in equation (12), becomes analytically challenging for a general transformation $g$. To overcome this difficulty, a common assumption in studies such as [25, 24] is that the low-fidelity function can be evaluated at zero cost. This eliminates the need for a surrogate for the low-fidelity function. Consequently, the non-linear transformation depicted in equation (13) transforms into creating a surrogate in a $d+1$ dimensional space. In our work, we delve into the modeling of $g$ using a Gaussian Process (GP). This approach enables us to effectively navigate the challenges associated with the joint distribution of surrogates, providing a viable solution for incorporating non-linear transformations within the multi-fidelity surrogate modeling framework.
@@
Figure 2 illustrates the relationship between $f_{l}$ and $f_{h}$. Modeling the transformation $g$ as described in equation (13) is synonymous with learning the manifold depicted in Figure 2. Crucially, this manifold lacks oscillations, signifying that learning it necessitates relatively fewer high-fidelity function evaluations. This observation underscores the advantage of multi-fidelity modeling in this example, as it allows for a learning process by leveraging the complicated feature to be captured in the low-fidelity function, thereby reducing the requirement of high-fidelity function evaluations.
@@
Here, $\theta_{\rho}$, $\theta_{f}$, and $\theta_{\delta}$ denote the kernel hyperparameters, which are determined by maximizing the likelihood, as discussed in the previous section. This kernel structure closely mirrors the Auto-Regressive Model (AR1) formulation described in equation (10), assigning each kernel to an individual part of the transformation. Specifically, $k_{\rho}$ models the scaling term, $k_{f}$ is responsible for transforming the low-fidelity model, and $k_{\delta}$ handles the additive correction term. The formulation above gives the method its name: Non-linear Auto-Regressive Gaussian Process (NARGP).
@@
Substituting equation (13) into equation (15), we obtain
@@
The Taylor expansion also motivates us to incorporate the derivative of low-fidelity in equation (13). Lee et al. [24] suggest writing the non-linear transformation as
@@
In many legacy simulators, the gradients are unavailable [39]. In cases where gradients are unavailable, one can resort to finite difference techniques to approximate the derivative. Lee et al. [24] suggest replacing the derivative in equation (16) with the delay term to accommodate such scenarios. This delay term evaluates the low-fidelity function after a small time step $\tau\in\mathbb{R}$. For a higher-dimensional input variable, we add a delay along each dimension, defining $\tau_{i}$ as a vector of zeros with $\tau$ at the $i^{\textrm{th}}$ position. The transformed formulation is now expressed as
@@
This method is called the Gaussian Process with Delay Fusion (GPDF) proposed in [24]. Additionally, we propose an enhancement to this method by modifying the kernel to mimic the Taylor expansion using a structured composite kernel similar to equation (14)
@@
Figure 3: Flowchart showing the training and prediction steps of NARGP surrogate for a three-fidelity scenario. $f_{1}$, $f_{2}$, and $f_{3}$ represent the function at each fidelity. $g_{1}$, $g_{2}$, and $g_{3}$ represent the surrogate of the corresponding fidelity. Training the NARGP model for a particular level involves evaluating the model of that fidelity level and the previous fidelity level. We draw samples to make predictions for a particular level, which involves drawing samples from the surrogate of the previous level and feeding them as input to the surrogate of that level. We can then use the samples to obtain the required statistical moments of the predictions by marginalizing the samples.
@@ 3.1.3 Deep Gaussian Processes
We model the transformation similar to equation (13) with one difference. We replace the actual function evaluation of the previous layer with the previous GP layer
@@
Figure 4: Flowchart showing the training and prediction steps of Deep Gaussian process surrogate for three fidelity scenarios. $f_{1}$, $f_{2}$, and $f_{3}$ represent the function at each fidelity. $g_{1}$, $g_{2}$, and $g_{3}$ represent the surrogate of the corresponding fidelity. Training the DGP model for a particular level involves evaluating the model of that fidelity level and the surrogate of the previous fidelity level, which further involves evaluating the one-level lower surrogate. This creates a cascading effect until we reach the lowest fidelity surrogate. The prediction process is similar to the full GP, which involves drawing and marginalizing samples.
@@
One can combine all the layers by multiplying the posterior from individual layers from equation (21) to obtain the joint posterior distribution for all layer $p(\mathbf{y},\{\mathbf{g}_{\ell}^{t},u_{\ell}\})$. Analytical inference from the posterior is intractable. To approximate the distribution, we assume that the prior $q(\mathbf{u}_{\ell})$ of $\mathbf{u}_{l}$ is a Gaussian distribution with mean $\tilde{\mathbf{m}}_{\ell}$ and variance $\tilde{\mathbf{S}}_{\ell}$. Now, the approximated DGP prior is
@@
If $\mathbf{u}_{\ell}$ is marginalized out of the equation (22), then the distribution becomes a Gaussian distribution with the following mean and variance:
@@
We can use the before approximating the posterior at each layer and then extend this to all the layers of DGP to obtain the approximate posterior distribution $q(\mathbf{y},\{\mathbf{g}_{\ell}^{\ell},\mathbf{u}_{\ell}\}_{\ell=1}^{L})$. We want to approximate the exact posterior distribution in equation (21) by an approximate distribution $q(\mathbf{y},\{\mathbf{g}_{\ell}^{\ell},\mathbf{u}_{\ell}\}_{\ell=1}^{L})$. We train the DGP by minimizing the Kullback-Leibler (KL) divergence $\textrm{KL}[q||p]$ between the variational posterior distribution $q$ and the true posterior distribution $p$. An exact evaluation of the KL divergence is not feasible. So, we convert the minimization of the KL divergence to the maximization of Evidence Lower Bound (ELBO) ($\mathcal{L}$) whose evaluation is feasible. ELBO is evaluated as
@@
The prediction process in Deep Gaussian Processes (DGP) follows a recursive transfer of results from the previous layer into the next layer. We draw samples at each layer from the approximate distribution defined by equation (22). We fuse the prediction parameter $X^{*}$ in intermediate layers with the sample drawn from the previous layer. We then use the fused parameter to draw a sample from that layer. We repeat this process until we have $S$ samples for each layer. We use the collected samples to compute the posterior mean and variance through a Monte Carlo approximation. Figure 4(b) visually illustrates these prediction steps for three fidelity models, depicting the recursive transfer and sampling process at each layer.
One can improve the surrogate by choosing a more structured kernel. Cutajar et al. [26] suggest using the kernel defined in equation (14). We call this method the Non-linear Auto-Regressive Deep Gaussian Process (NARDGP). We suggest using the delay term to further add useful information. We call the method Deep Gaussian Process with Delay Fusion (DGPDF). We further improve upon the suggested method using a structured kernel as mentioned in equation (18). We named the resulting method Deep Gaussian Process with Delay Fusion and Composite kernel (DGPDFC).
@@ 3.2 Adaptive model improvement
Upon simplification, it can be shown that for GPs, the gain in information is directly proportional to the logarithm of the posterior predicted standard deviation. Thus, the optimization problem stated in equation (27) can be converted to
@@
There are certain drawbacks to this method. Firstly, this method adds one point at a time, limiting its overall speed. Secondly, the suggested points are always towards the boundary in high-dimensional space.
Various other approaches are proposed in the literature depending on alternative metrics to iteratively enhance the Gaussian Process surrogate [48]. One can improve the posterior standard deviation adaptivity metric by adjusting it via some error information as discussed in [49]. However, this requires function evaluation at additional validation points, making it infeasible for computationally expensive high-fidelity functions. Query-by-committee (QBC) [50, 51] is also computationally infeasible due to the need to train multiple models. Gradient-based adaptive approach [52] utilizes gradient information, which may be unavailable in many cases especially involving legacy solvers. Moreover, gradient approximation using the finite difference method is numerically unstable and computationally expensive for high-fidelity models [53]. There are still many other adaptivity algorithms [48, 54, 55]. We will leave the discussion and the effects of those algorithms on multi-fidelity GP as future works.
@@ 4 Numerical Examples
In this segment, we evaluate the surrogate modeling techniques mentioned in Section 3 using standard academic examples and two real-world problems, each presenting its distinct challenges. Subsequently, we conduct a comparative and critical analysis of the strengths and limitations of each approach in varied scenarios. The method outlined in Section 3 was implemented in Python programming language using GPflow library [56]. This implementation is open source and can be accessed on GitHub [^1] https://github.com/KislayaRavi/MuDaFuGP.